# Safe Reinforcement Learning In Black-Box Environments Via Adaptive Shielding

### v2

@@ Abstract
Empowering safe exploration of reinforcement learning (RL) agents during training is a critical challenge towards their deployment in many real-world scenarios. When prior knowledge of the domain or task is unavailable, training RL agents in unknown, black-box environments presents an even greater safety risk. We introduce ADVICE (Adaptive Shielding with a Contrastive Autoencoder), a novel post-shielding technique that distinguishes safe and unsafe features of state-action pairs during training, and uses this knowledge to protect the RL agent from executing actions that yield likely hazardous outcomes. Our comprehensive experimental evaluation against state-of-the-art safe RL exploration techniques shows that ADVICE significantly reduces safety violations ($\approx\!\!50\%$) during training, with a competitive outcome reward compared to other techniques.
Safe Reinforcement Learning, Reinforcement Learning, Safe Exploration
@@ 1 Introduction
Reinforcement Learning (RL) (Sutton and Barto, 2018) is a powerful machine learning paradigm for solving complex decision-making tasks. RL has exhibited performance commensurate with the cognitive abilities of humans in diverse applications, including game-playing (Berner et al., 2019) and robot control (Rudin et al., 2022; Heess et al., 2017). Despite its huge potential, developing RL-based agents that can explore their environment safely remains a major challenge. Exploring unfamiliar, and potentially hazardous, states while learning from the environment, poses real dangers especially in safety-critical domains, like robotics or healthcare. Alleviating this requires RL agents that can synthesise an optimal policy by exploring the policy space adequately while ensuring safe exploration by avoiding unsafe actions (Amodei et al., 2016).
@@
Ensuring safety becomes an increasingly difficult challenge as environments characterised by continuous state/action spaces become increasingly complex (Dalal et al., 2018). Such environments require a large amount of training time before the agent can consistently complete the task and avoid safety concerns. This issue is further exacerbated in black-box environments where no prior knowledge can be utilised before training; only data observed in real-time by the RL agent is available. In such scenarios, the risk associated with exploration increases as the agent must operate without pre-defined guidelines, rendering typical safe exploration techniques inadequate (Waga et al., 2022).
@@
Prior research on safe RL exploration formulates safety constraints as linear temporal (Könighofer et al., 2023) and probabilistic logic (Yang et al., 2023) specifications, whose use as a shield protects the agent during training. Shielding techniques are categorised into pre-shielding (restricting action choices to a predefined safe subset) and post-shielding (evaluating and modifying actions post-selection to ensure safety) (Odriozola-Olalde et al., 2023). Despite major advances, a common challenge across these methods is their reliance on some degree of prior knowledge about the environment, task, or safety concern—a strong assumption that is not generally applicable. Research targeting safe exploration in black-box environments employs Lagrangian methods (Stooke et al., 2020; Achiam et al., 2017), or involves a pre-training phase before the shield synthesis (Tappler et al., 2022). However, these methods require an explicit cost signal at every timestep or a small amount of prior knowledge to initialize the shield, neither of which is available in many environments.
@@
Figure 1: High-level ADVICE overview, showing its shield construction (A), and execution and adaptation (B) stages. During shield construction, safe and unsafe features (signifying accepting and non-accepting terminal states, respectively) are collected and used by a contrastive autoencoder to learn latent representations (embeddings). During shield execution and adaptation, a new feature is encoded into the contrastive latent space and the nearest neighbours guide its classification as safe or unsafe. ADVICE’s cautiousness is adapted dynamically based on the RL agent’s recent safety performance.
@@
We mitigate this significant limitation by introducing ADVICE (ADaptiVe ShIelding with a Contrastive AutoEncoder), a novel post-shielding technique for safe RL-agent exploration in black-box environments. Figure 1 shows a high-level overview of ADVICE’s two-stage approach. During shield construction (Figure 1-A), ADVICE uses a contrastive autoencoder to learn latent representations that distinguish safe and unsafe features (i.e., state-action pairs). During execution (Figure 1-B), ADVICE employs the learnt representations and a non-parametric classifier to identify and correct unsafe actions. It also adapts its cautiousness based on the agent’s recent performance, encouraging exploration when suitable. ADVICE demonstrates effectiveness in complex black-box environments, significantly reducing safety violations compared to state-of-the-art methods like Lagrangian (Lillicrap et al., 2015) and discretized shields (Shperberg et al., 2022), and conservative safety critics (Bharadhwaj et al., 2020). To the best of our knowledge, ADVICE is the first shielding approach for safe RL in black-box environments with continuous state/action spaces, offering an end-to-end shield synthesis without using any prior knowledge.
@@ 2 Related Work
Shielding Techniques. Recent research has devised techniques for the safe exploration of RL agents using safety shields, allowing the agent to select from a pool of safe actions or correct an action deemed unsafe (Odriozola-Olalde et al., 2023). Existing shielding approaches leverage linear temporal logic (LTL) specifications (Alshiekh et al., 2018; Könighofer et al., 2023; ElSayed-Aly et al., 2021) or use external hints for constructing the LTL formulae (Waga et al., 2022). LTL specifications can be replaced with probabilistic logic programming (Yang et al., 2023), extending their use to continuous deep RL and enabling safety constraints to be differentiable. By utilising logical neural networks (Kimura et al., 2020), the same logic specifications can be both respected and learnt, providing a more nuanced understanding of safety. Jansen et al. (2020) introduce probabilistic shields to ensure safety, while Carr et al. (2023) implement safety under partial observability and Goodall and Belardinelli (2023) use approximate models of the environment to maintain safety. Dalal et al. (2018) train a safety layer to adhere to a known constraint function. Although these methods can improve safety and, in limited cases, eliminate violations, their demand for explicit prior knowledge of the RL agent environment, task, and/or safety concerns, restricts their applicability to a narrow set of environments and safety aspects (Turchetta et al., 2020). In contrast, our ADVICE post-shielding method does not need any prior knowledge, and uses exclusively the information captured in a typical RL problem.
@@
Black-Box Safe Exploration Techniques. Other recent research focuses on improving safety in black-box environments, where no prior knowledge is provided to the agent/user. A trivial but effective solution is to record all unsafe features in a tabular format to prevent the agent from repeating them (Shperberg et al., 2022). However, the applicability of this approach is limited to discrete environments or extremely low-dimensional spaces. Other solutions collect environment data before training, to then instrument a safety layer (Srinivasan et al., 2020; Thananjeyan et al., 2021; Bharadhwaj et al., 2020) or shield (Tappler et al., 2022) that protects the agent. This requires a significant amount of data collection, resulting in a prohibitive number of safety violations. Lagrangian methods enable modelling the Markov Decision Process encoding the RL problem as a Constrained Markov Decision Process, and are widely adopted due to their simplicity and effectiveness (Altman, 2021; Garcıa and Fernández, 2015). The Lagrangian multiplier can be fixed (Stooke et al., 2020; Altman, 1998), or integrated into the algorithm itself (Tessler et al., 2018; Achiam et al., 2017). Other solutions employ uncertainty estimation concepts (Kahn et al., 2017; Jain et al., 2021). Defining safety in terms of uncertainty and propagating it into the RL algorithm during training yields a cautious yet effective agent for reducing safety violations, even in complex environments. Similarly to these approaches, ADVICE needs no prior knowledge, not even a cost signal; however, ADVICE also needs far less data collection than the previously discussed solutions. For example, in Tappler et al. (2022), the shielding method needs approximately $10,000$ episodes to begin promoting safety successfully, whereas we show ADVICE promoting safety by using only between $500$ to $1,000$ episodes in Figure 12.
@@ 3 Preliminaries
Markov Decision Process. A Markov Decision Process (MDP) (Bellman, 1957) is a discrete-time stochastic control process to model decision-making. An MDP is formally defined as a 5-tuple $M=(S,A,P,R,\gamma)$, where $S$ is the state space, $A$ is the action space, $P$ is the state transition probability matrix/transition function such that $P(s^{\prime}|s,a)$ is the probability of transitioning to state $s^{\prime}$ from state $s$ using action $a$, $R$ is the reward function such that $R(s,a)$ is the reward for taking action $a$ in state $s$, and $\gamma$ is the discount factor that determines the value of future rewards. A policy $\pi:S\to\Delta(A)$ is a distribution over actions given a state. MDPs can be solved using dynamic programming techniques (e.g., value iteration, policy iteration) which require complete knowledge of the MDP’s dynamics (Bertsekas and Tsitsiklis, 2008). An MDP includes accepting states, which are typically associated with the partial or full completion of a task, as well as non-accepting terminal states, which signify the end of the process and are often used to represent safety violations.
[…]
Contrastive Learning. Contrastive Learning (CL) (Hadsell et al., 2006) is an unsupervised or semi-supervised machine learning paradigm aimed at distinguishing between similar (positive) and dissimilar (negative) pairs of data points. At its core lies a contrastive loss function, which encourages the model to put similar pairs closer together in the embedding space while separating dissimilar pairs. Given a pair of inputs $x_{i}$, $x_{j}$, the contrastive loss function is defined as:
@@ 4 ADVICE
The ADVICE post-shielding technique empowers safe RL exploration in black-box environments without requiring a system model. The core of ADVICE is a contrastive autoencoder (CA) model that can efficiently distinguish between safe and unsafe features from the feature space $\mathcal{F}:S\times A$ (representing all possible state-action pairs). A feature $f_{t}=(s_{t},a_{t})$, $f_{t}\in\mathcal{F}$, denotes the state-action pair at timestep $t$. The CA model leverages a unique loss function where similar and dissimilar features are compared, enabling the systematic identification of meaningful latent feature representations. ADVICE employs these latent representations and trains an unsupervised nearest neighbours model to the learnt embedding space, thus enabling the classification of new features. Formally, the ADVICE shield is defined by a function $\phi:F\times\mathbb{Z}_{+}\rightarrow A$, such that at time step $t$ during the execution of an episode, ADVICE evaluates the agent’s desired action $a_{t}$ by computing $\phi(f_{t},K)$, where $K$ is a parameter that controls the risk aversion level of the shield. This function returns $a_{t}$ if the desired action is deemed safe (thus endorsing its execution), or returns (and enforces the execution of) another, safe action $a^{\prime}_{t}$ instead. ADVICE considers the agent’s performance and automatically adapts the value of $K$, thus supporting the dynamic calibration of ADVICE’s cautiousness level during learning. We detail next the two ADVICE stages, and its use within the RL loop as a post-shield.
@@ 4.1 ADVICE Shield Construction
As a black-box post-shield, ADVICE does not rely on any prior knowledge about the RL agent or its environment. Instead, the shield construction stage is founded on collecting a feature set $F_{E}$ during an initial, unshielded interaction period of $E$ episodes where the RL agent is allowed to interact with and collect experience from its environment. A feature $f_{t}\in F_{E}$ is then classified as safe, unsafe or inconclusive based on the following:
$g(f_{t})\!\!=\!\!\begin{cases}\text{safe},&\text{if }t=0\text{ or }s_{t+1}\text{ is an accepting state}\\ \text{unsafe},&\text{if }s_{t+1}\text{ is a non-accepting terminal state}\\ \text{inconc},&\text{otherwise}\end{cases}$ (2)
and the set $F_{E}$ is partitioned into disjoint subsets of safe, unsafe and inconclusive features: $S=\{f_{t}\!\in\!F_{E}|g(f_{t})=\text{safe}\}$, $U\!=\!\{f_{t}\!\in\!F_{E}|g(f_{t})=\text{unsafe}\}$, and $I\!=\!F_{E}\setminus(S\cup U)$. This partitioning facilitates the contrastive learning of: (i) similar feature pairs (i.e., two safe or two unsafe features); and (ii) dissimilar feature pairs (i.e., a safe and an unsafe feature), allowing the CA model to discern between safe and unsafe features effectively, and to focus on finding meaningful representations in a lower-dimensional latent space that adequately reflects the problem domain similarities and differences. We assemble a training dataset $\mathcal{C}$ for the CA model by labelling each combination of distinct features from $S\cup U$ as similar (label ‘1’) or dissimilar (‘0’):
$\mathcal{C}=$$\{((f_{t},f_{t^{\prime}}),1)\mid f_{t},f_{t^{\prime}}\in S\cup U\wedge$ (3)
@@
$\{((f_{t},f_{t^{\prime}}),0)\mid f_{t},f_{t^{\prime}}\in S\cup U\wedge g(f_{t})\neq g(f_{t^{\prime}})\}$
Using this dataset, the CA model is trained to optimise two loss functions: mean squared error (MSE) and contrastive loss (CL) from Equation (1). The MSE measures how accurately the model reconstructs input features, while the CL refines the latent space by clustering similar features and separating dissimilar ones based on Euclidean distance. This encourages cohesion among similar features and separation among dissimilar ones. As such, the ADVICE CA model learns to encode and reconstruct the salient features of the RL problem – environment state and chosen action – accurately in a lower-dimensional latent space.
Once trained, the CA model is adept at finding nuanced distinctions between safe and unsafe features, and accurately placing unseen features within the appropriate partitions in the latent space. The shield construction stage of ADVICE concludes with embedding an unsupervised nearest neighbours (KNN) model in the latent space; this model classifies new encoded features as safe/unsafe. A visual representation of the latent data encoding, illustrating the clear separation in the latent space, is shown in Figure 1.
@@
[…]
@@
2:     $s_{1}\leftarrow\textsc{Observe()}$(), $t\leftarrow 1$
@@
22:     $K\leftarrow\textsc{UpdateK}(K,K^{max},(z_{i})_{i=1}^{e},h_{d},h_{r})$
@@
[…]
27:   $N_{\hat{f}}\leftarrow\textsc{GetSafeNeighbours}(\hat{f},K^{max})$

### v3

@@ Abstract
Safe exploration of reinforcement learning (RL) agents is a critical activity for empowering their deployment in many real-world scenarios. When prior knowledge of the target domain or task is unavailable, training RL agents in unknown, black-box environments unavoidably yields significant safety risks. Our ADVICE (Adaptive Shielding with a Contrastive Autoencoder) novel post-shielding approach operates in continuous state and action spaces, distinguishing safe and unsafe features of state-action pairs during training, and uses this knowledge to safeguard the RL agent from executing actions that yield likely hazardous outcomes. Our comprehensive experimental evaluation shows that ADVICE significantly reduces safety violations ($\approx\!\!50\%$) compared to state-of-the-art safe RL exploration approaches, while maintaining a competitive outcome reward for the synthesised safe policy.
@@ 1 Introduction
Reinforcement Learning (RL) [37] is a powerful machine learning paradigm for solving complex decision-making tasks. RL has exhibited performance commensurate with the cognitive abilities of humans in diverse applications, including game-playing [9] and robot control [32, 21]. Despite its huge potential, developing RL-based agents that can explore their environment safely remains a major challenge. Exploring unfamiliar, and potentially hazardous, states while learning from the environment, poses real dangers especially in safety-critical domains, like robotics or healthcare. Alleviating this challenge requires RL agents that can safely explore their environment by avoiding unsafe actions while simultaneously devising an optimal policy by exploring the policy space adequately [6].
@@
Ensuring safety becomes an increasingly difficult challenge as environments characterised by continuous state/action spaces become progressively more complex [16]. Such environments require a large amount of training time before the RL agent can consistently complete the task and avoid safety concerns. This issue is further exacerbated in black-box environments where prior or domain knowledge is not available in advance; only data observed in real-time by the RL agent is available. In such scenarios, the risk associated with exploration increases exponentially as the agent must operate without pre-defined guidelines or guardrails, rendering typical safe exploration approaches inadequate [41].
@@
Prior research on safe RL exploration formulates safety constraints as linear temporal [28] and probabilistic logic [43] specifications, whose use as a shield protects the agent during training. Shielding approaches are categorised into pre-shielding (restricting action choices to a predefined safe subset) and post-shielding (evaluating and modifying actions post-selection to ensure safety) [31]. Despite major advances, a common challenge across these approaches is their reliance on some degree of prior knowledge about the environment, task, or safety concern—a strong assumption that is not generally applicable [41]. Research targeting safe exploration in black-box environments employs Lagrangian approaches [2, 36], or involves a pre-training phase before the shield synthesis [38]. However, these approaches require an explicit cost signal at every timestep or a small amount of prior knowledge to initialize the shield, neither of which is available in many environments.
@@
Figure 1: High-level ADVICE overview, showing its shield construction (A), and execution and adaptation (B) stages. During shield construction, safe and unsafe features (signifying accepting and non-accepting terminal states, respectively) are collected and used by a contrastive autoencoder to learn latent representations (embeddings). During shield execution and adaptation, a new feature is encoded into the contrastive latent space and the nearest neighbours guide its classification as safe or unsafe. ADVICE’s cautiousness is adapted dynamically based on the RL agent’s recent safety performance. A detailed ADVICE description is provided in Section 4.
@@
We mitigate this significant limitation by introducing ADVICE (ADaptiVe ShIelding with a Contrastive AutoEncoder), a novel post-shielding approach for safe RL agent exploration in black-box environments with continuous state and action spaces, where no prior knowledge of the environment dynamics, and reward structure is available. Following [3, 7], ADVICE relaxes traditional cost-based safety assumptions by operating within the standard MDP framework, defining safety through accepting and non-accepting terminal states, thereby enabling greater generalisability across tasks and environments without requiring explicit cost signals. Unlike many prior shielding approaches, ADVICE does not rely on handcrafted domain knowledge or expert-defined safety constraints, enabling full end-to-end shield synthesis in black-box environments. To the best of our knowledge, ADVICE is the first shielding approach for safe RL in black-box environments with continuous state/action spaces, offering an end-to-end shield synthesis without using any prior knowledge. Figure 1 shows a high-level overview of ADVICE’s two-stage approach. During shield construction (Figure 1A), ADVICE uses a contrastive autoencoder to learn latent representations that distinguish safe and unsafe features (i.e., state-action pairs). During execution (Figure 1B), ADVICE employs the learnt representations and a non-parametric classifier to identify and correct unsafe actions. ADVICE also dynamically adapts its cautiousness based on the agent’s recent performance, encouraging exploration when suitable. Finally, ADVICE demonstrates effectiveness in complex black-box environments, significantly reducing safety violations compared to state-of-the-art approaches like Lagrangian [29] and discretized shields [34], and conservative safety critics [13].
@@ 2 Related Work
Shielding Approaches. Recent research has devised approaches for the safe exploration of RL agents using safety shields, allowing the agent to select from a pool of safe actions or correct an action deemed unsafe [31]. Existing shielding approaches leverage linear temporal logic (LTL) specifications [3, 17, 28] or use external hints for constructing the LTL formulae [41]. LTL specifications can be replaced with probabilistic logic programming [43], extending their use to continuous deep RL and enabling safety constraints to be differentiable. By utilising logical neural networks [27], the same logic specifications can be both respected and learnt, providing a more nuanced understanding of safety. Jansen et al. [23] introduce probabilistic shields to ensure safety, while Carr et al. [15] implement safety under partial observability and Goodall and Belardinelli [19] use approximate models of the environment to maintain safety. Dalal et al. [16] train a safety layer to adhere to a known constraint function. Although these approaches can improve safety and, in limited cases, eliminate violations, their demand for explicit prior knowledge of the RL agent environment, task, and/or safety concerns, restricts their applicability to a narrow set of environments and safety aspects [40]. In contrast, our ADVICE post-shielding approach does not need any handcrafted prior knowledge, and uses exclusively the information captured in a typical RL problem.
@@
Black-Box Safe Exploration Approaches. Other recent research focuses on improving safety in black-box environments, where no prior knowledge is provided to the agent/user. A trivial but effective solution is to record all unsafe features in a tabular format to prevent the agent from repeating them [34]. However, the applicability of this approach is limited to discrete environments or extremely low-dimensional spaces. Other solutions collect environment data before training, to then instrument a safety layer [35, 13] or shield [38] that protects the agent. This requires a significant amount of data collection, resulting in a prohibitive number of safety violations. Lagrangian approaches enable modelling the Markov Decision Process, encoding the RL problem as a Constrained Markov Decision Process, and are widely adopted due to their simplicity and effectiveness [5, 18]. The Lagrangian multiplier can be fixed [4, 36], or integrated into the algorithm itself [2, 39]. Other solutions employ uncertainty estimation concepts [22, 26]. Defining safety in terms of uncertainty and propagating it into the RL algorithm during training yields a cautious yet effective agent for reducing safety violations, even in complex environments. Similarly to these approaches, ADVICE needs no prior knowledge, not even a cost signal;ADVICE also needs far less data collection than the previously discussed solutions. For example, in Tappler et al. [38], the shielding approach needs approximately $10,000$ episodes to begin promoting safety successfully, whereas we show ADVICE starts safeguarding the RL agent by using only between $500$ to $1,000$ episodes in Figure 14.
@@ 3 Preliminaries
Markov Decision Process. A Markov Decision Process (MDP) [8] is a discrete-time stochastic control process to model decision-making. An MDP is formally defined as a 5-tuple $M=(S,A,P,R,\gamma)$, where $S$ is the state space, $s_{t}$ is the state at timestep $t$, AA is the action space, $P$ is the state transition probability matrix/transition function such that $P(s_{t+1}|s_{t},a_{t})$ is the probability of transitioning to state $s_{t+1}$ from state $s_{t}$ using action $a_{t}$, $R$ is the reward function such that $R(s,a)$ is the reward for taking action $a$ in state $s$, and $\gamma$ is the discount factor that determines the weight of future rewards. A policy $\pi:S\to\Delta(A)$ is a distribution over actions given a state. MDPs can be solved using dynamic programming techniques (e.g., value iteration, policy iteration), which require complete knowledge of the MDP’s dynamics [11]. An MDP includes accepting states, which are typically associated with the partial or full completion of a task, as well as non-accepting terminal states, which signify the end of the process and are often used to represent safety violations. In this work, we consider MDPs with continuous state and action spaces. That is, $S\subseteq\mathbb{R}^{n}$ and $A\subseteq\mathbb{R}^{m}$ are real and possibly high-dimensional values.
Reinforcement Learning. Reinforcement Learning (RL) involves training an agent to make a sequence of decisions by interacting with the environment, which is modelled as a Markov Decision Process (MDP). This machine learning technique solves MDPs when full knowledge of the transition function is not available. The agent’s goal is to find a policy $\pi^{*}$ maximising the expected discounted return $E\left[\sum_{t=0}^{\infty}\gamma^{t}R(s_{t},a_{t})\right]$ [37]. A value function $V_{\pi}(s)$ informs the agent how valuable a given state is when following the current policy $\pi$. Regret, a fundamental concept in RL [10], measures the difference between the cumulative reward of an optimal policy and the cumulative reward of the policy followed by the agent, highlighting the cost of learning and exploration over time. The standard regret factor of the RL process, which reflects the optimal trade-off between exploration and exploitation, is $O(\frac{1}{\sqrt{E_{\mathit{max}}}})$, where $E_{\mathit{max}}$ is the maximum number of episodes the agent is trained for. Common RL algorithms include Q-learning [42], and SARSA [33]. Deep reinforcement learning (DRL) extends traditional algorithms by utilising deep neural networks to approximate the policy $\pi$ or the value function $V$ when the state/action space is high-dimensional and complex. Actor-critic approaches [37] are a popular class of algorithms both in traditional and deep RL. Distinctly, the policy (actor) and the value function (critic) are modelled as separate components, allowing for simultaneous updates to both functions. Deep Deterministic Policy Gradient (DDPG) [29] is an example of an actor-critic approach tailored specifically for continuous action spaces.
Contrastive Learning. Contrastive Learning (CL) [20] is an unsupervised or semi-supervised machine learning paradigm aimed at distinguishing between similar (positive) and dissimilar (negative) pairs of data points. At its core lies a contrastive loss function, which encourages the model to put similar pairs closer together in the embedding space while separating dissimilar pairs. Given a pair of inputs $x_{i}$, $x_{j}$, the contrastive loss function is defined as:
@@ 4 ADVICE
The ADVICE post-shielding approach  [^1] Code available at: https://github.com/team-daniel/ADVICE (Figure 1) empowers safe RL exploration in black-box environments without requiring a system model, domain knowledge or handcrafted rules as in related research [28, 43]. The core of ADVICE is a contrastive autoencoder (CA) model that can efficiently distinguish between safe and unsafe features from the feature space $F:S\times A$ (representing all possible state-action pairs). A feature $f_{t}=(s_{t},a_{t})$, $f_{t}\in F$, denotes the state-action pair at timestep $t$. The CA model leverages a unique loss function where similar and dissimilar features are compared, enabling the systematic identification of meaningful latent feature representations. ADVICE employs these latent representations and trains an unsupervised nearest neighbours model in the learnt embedding space, thus enabling the classification of new features. Formally, the devised ADVICE shield is defined by a function $\phi:F\times\mathbb{Z}_{+}\rightarrow A$, such that at time step $t$ during the execution of an episode, ADVICE evaluates the agent’s desired action $a_{t}$ by computing $\phi(f_{t},K)$, where $K$ is a parameter that controls the risk aversion level of the shield. This function returns $a_{t}$ if the desired action is deemed safe (thus endorsing its execution), or returns (and enforces the execution of) another, safe action $a^{\prime}_{t}$ instead. ADVICE considers the agent’s performance and automatically adapts the value of $K$, thus supporting the dynamic calibration of ADVICE’s cautiousness level during learning. We detail next the two ADVICE stages, and their use within the RL loop as a post-shield.
@@ 4.1 ADVICE Shield Construction
As a black-box post-shield, ADVICE does not rely on any domain knowledge about the RL agent or its environment. Instead, the shield construction stage is founded on collecting a feature set $F_{E}$ during an initial, unshielded interaction period of $E$ episodes, where the RL agent is allowed to interact with and collect experience from its environment. This relaxes the safety definition from assuming a cost-based MDP framework (like the approaches from [13, 14, 39]) to the standard MDP (as in [3, 7]), thereby enabling greater generalisability. A feature $f_{t}\in F_{E}$ collected up to episode $E$ is then classified as safe, unsafe or inconclusive based on the following:
[…]
and the set $F_{E}$ is partitioned into disjoint subsets of safe, unsafe and inconclusive features: $F_{safe}=\{f_{t}\!\in\!F_{E}|g(f_{t})=\text{safe}\}$, $F_{unsafe}\!=\!\{f_{t}\!\in\!F_{E}|g(f_{t})=\text{unsafe}\}$, and $F_{inc}\!=\!F_{E}\setminus(F_{safe}\cup F_{unsafe})$. This partitioning facilitates the contrastive learning of: (i) similar feature pairs (i.e., two safe or two unsafe features); and (ii) dissimilar feature pairs (i.e., a safe and an unsafe feature), allowing the CA model to discern between safe and unsafe features effectively. This key ADVICE characteristic allows to focus on finding meaningful representations in a lower-dimensional latent space that adequately reflects the problem domain similarities and differences. We assemble a training dataset $\mathcal{C}$ for the CA model by automatically labelling each combination of distinct features from $F_{safe}\cup F_{unsafe}$ as similar (label ‘1’) or dissimilar (‘0’):
$\mathcal{C}=$ $\{((f_{t},f_{t^{\prime}}),1)\mid f_{t},f_{t^{\prime}}\in F_{safe}\cup F_{unsafe}\wedge$ (3)
@@
$\{((f_{t},f_{t^{\prime}}),0)\mid f_{t},f_{t^{\prime}}\in F_{safe}\cup F_{unsafe}\wedge g(f_{t})\neq g(f_{t^{\prime}})\}$
Using this dataset, the CA model is trained to optimise two loss functions: contrastive loss (CL) from Equation (1) and mean squared error (MSE). The MSE measures how accurately the model reconstructs input features, while the CL refines the latent space by clustering similar features and separating dissimilar ones based on Euclidean distance. This encourages cohesion among similar features and separation among dissimilar ones. Consequently, the ADVICE CA model learns to encode and reconstruct the salient features of the RL problem – environment state and chosen action – accurately in a lower-dimensional latent space.
Once trained, the CA model is adept at finding nuanced distinctions between safe and unsafe features, and accurately placing unseen features within the appropriate partitions in the latent space. The shield construction stage of ADVICE concludes with embedding an unsupervised nearest neighbours (KNN) model in the latent space; this model classifies new encoded features as safe/unsafe. A visual representation of the latent data encoding, illustrating the clear separation in the latent (embedding) space, is shown in Figure 1.
@@
0: Contrastive autoencoder $\mathit{CA}$, Max neighbours $K_{max}$, Safety threshold $K$, Safety violation indicators $(z_{i})_{i=1}^{E}$, Recent/distant time windows $h_{r}$, $h_{d}$, Shield construction episodes $E$, Max episodes $E_{max}$, Max timesteps $T$, Policy $\pi$
@@
2:  $s_{1}\leftarrow\textsc{Observe()}$, $t\leftarrow 1$
@@
22:  $K\leftarrow\textsc{UpdateK}(K,K_{max},(z_{i})_{i=1}^{e},h_{d},h_{r})$
@@
25:function IsSafe($f,K,K_{max},\mathit{CA}$) )
26:  $\hat{f}\leftarrow\textsc{Encode}(f,\mathit{CA})$
27:  $N_{\hat{f}}\leftarrow\textsc{GetSafeNeighbours}(\hat{f},K_{max})$